OpenStudy (anonymous):

Task 1 Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so is \[f(x) = \frac{x + a}{b}\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok so look at the inverse... swap x and y \[x = \frac{y + a}{b}\] can you make y the subject...?

OpenStudy (anonymous):

no?

OpenStudy (campbell_st):

ok... multiply both sides by b \[bx = y + a\] now you should be able to make y the subject...

OpenStudy (anonymous):

how?

OpenStudy (campbell_st):

subtract a from both sides...

OpenStudy (anonymous):

\[a + bx = y\]

OpenStudy (campbell_st):

well its subtract.. y = bx - a now compare it to the inverse you are given y = cx - d what can you say about b, c and a and d..?

OpenStudy (anonymous):

they are both the same ?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

So, when you do the inverse, it like the same?

OpenStudy (campbell_st):

it requires you to swap x and y and then make y the subject... so all you need to do, from the way I see it... is to find the inverse of f(x) compare it to g(x) and make the conclusion c = b and a = d then pick numbers for a and b using the conclusion you will also have c and d

OpenStudy (anonymous):

ohh ok